### v1

# On the Genericity of the Spectrum Intervalization for Multi-Frequency Quasiperiodic Schrödinger Operators

Abstract
This paper proves a genericity conjecture by Goldstein, Schlag, and Voda[Invent. Math.217(2019)] for multi-frequency quasiperiodic Schrödinger operators. Specifically, we show that for almost all coefficients of real trigonometric polynomial potentials, the spectrum forms a single interval under strong coupling conditions. This confirms a long-standing intuition by Chulaevky and Sinai[Comm.Math.Phys.125(1989)] that the spectrum typically intervals for generic potentials, and extends the existence results of Goldstein et al. to a full measure setting. Our proof relies on tools from differential topology, measure theory, and analytic function theory.
Mathematics Subject Classification (2020): 47A10, 47B39
keywords:
Quasiperiodic Schrödinger operators , Spectrum , Genericity , Cartan estimates , Transversality theory
journal: arXiv
1 Introduction
Quasiperiodic Schrödinger operators have been extensively studied in mathematical physics, particularly in the context of Anderson localization [4, 3, 6] and spectral theory [5, 9]. A fundamental question concerns the structure of the spectrum: whether it is a Cantor set [1] or a single interval [11]. For multi-frequency operators with analytic potentials, it was first suggested by Chulaevsky and Sinai [8] that under strong coupling, the spectrum typically forms an interval for generic potentials. This intuition was later formalized by Goldstein, Schlag, and Voda [11], who proved that for a specific class of potentials (denoted class $\mathfrak{G}$), the spectrum is indeed an interval. Moreover, they conjectured that class $\mathfrak{G}$ is generic, i.e., holds for almost all coefficients in the space of trigonometric polynomials.
In this paper, we prove the genericity conjecture by Goldstein et al. [11], showing that for almost all coefficients of real trigonometric polynomials, the potential belongs to class $\mathfrak{G}$, which ensures the spectrum is a single interval. Our approach combines tools from differential topology, measure theory, and analytic function theory, specifically leveraging the parametric transversality theorem, and Cartan-type estimates to establish the full-measure property of class $\mathfrak{G}$. This provides a comprehensive framework for understanding the generic behavior of these operators and confirms the original intuition of Chulaevsky and Sinai [8].
2 Preliminaries
2.1 Class $\mathfrak{G}$ of Potetials
Consider the multi-frequency quasiperiodic Schrödinger operator on $\ell^{2}(\mathbb{Z})$:
$(H(x)\psi)(n)=-\psi(n+1)-\psi(n-1)+\lambda V(x+n\omega)\psi(n),$ (2.1)
where $x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}$ is a Diophantine frequency vector, $\lambda>0$ is the coupling constant, and $V$ is a real analytic potential.
We focus on trigonometric polynomial potentials of the form:
$V(x)=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x},$ (2.2)
with coefficients $\mathbf{c}=(c_{m})\in\mathbb{R}^{N}$, where $N$ is the number of integer vectors $m\in\mathbb{Z}^{d}$ satisfying the condition $|m|=\sum_{j=1}^{d}|m_{j}|\leq n$. By a combinatorial count, the dimension $N$ is given explicitly by:
$N=\sum_{k=0}^{n}a_{d,k},\quad\text{where}\quad a_{d,k}=\sum_{l=1}^{\min(d,k)}\binom{d}{l}2^{l}\binom{k-1}{l-1}\ \text{for}\ k\geq 1,\ \text{and}\ a_{d,0}=1.$ (2.3)
This counts the number of solutions to the inequality $|m|\leq n$ in integers.
Goldstein et al. [11] introduced the following function class:
Definition 2.1 (Class $\mathfrak{G}$).
A potential $V$ belongs to class $\mathfrak{G}$ if it satisfies the following conditions:
  1. (i)
    $V$ is a Morse function (all critical points are non-degenerate).
  2. (ii)
    $V$ has unique global minimum and maximum points.
  3. (iii)
    For any $h\in\mathbb{T}^{d}$ with $\|h\|\geq\exp(-K)$ and $K\gg 1$, the Cartan estimate holds:
    ${\rm mes}\left\{x\in\mathbb{T}^{d}:\min\left(|V(x+h)-V(x)|,|g_{V,h,i,j}(x)|\right)<\exp(-K)\right\}$
    $\quad\leq\exp(-K^{\mathfrak{c}_{1}}),$ (2.4)
    where
    $g_{V,h,i,j}(x)=\operatorname{det}\left[\begin{array}[]{cc}\partial_{x_{i}}V(x)&\partial_{x_{j}}V(x)\\ \partial_{x_{i}}V(x+h)&\partial_{x_{j}}V(x+h)\end{array}\right].$
    (see [7] for the origin of such estimates).
  4. (iv)
    For any unit vector $h_{0}\in\mathbb{R}^{d}$ and $\eta\in\mathbb{R}$, the gradient Cartan estimate holds:
    ${\rm mes}\left\{x\in\mathbb{T}^{d}:\min\left(|V(x)-\eta|,|\langle\nabla V(x),h_{0}\rangle|\right)<\exp(-K)\right\}\leq\exp(-K^{\mathfrak{c}_{1}}).$ (2.5)
Here, $\mathfrak{c}_{1}>0$ is a constant, and the Cartan estimates are based on the theory of analytic functions as developed by Cartan [7] and applied in [11].
2.2 Statement of Main Theorem
We now state the main theorem, which proves the genericity conjecture by Goldstein et al. [11, Remark 1.2(b)].
Theorem 2.2 (Genericity of Class $\mathfrak{G}$).
Let $V$ be a trigonometric polynomial potential as in (2.2). Then the set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which $V$ belongs to class $\mathfrak{G}$ has full Lebesgue measure. Consequently, for almost all $\mathbf{c}$, under strong coupling ($\lambda\gg 1$), the spectrum of the operator $H(x)$ is a single interval.
The proof of Theorem 2.2 is based on the parametric transversality theorem from differential topology.
2.3 Parametric Transversality Theorem
We now recall the essential tools from differential topology [2, 12, 13, 14] which are used to prove that key properties hold for almost every potential in our genericity analysis. These results are foundational for establishing that exceptional sets have measure zero in finite-dimensional parameter spaces.
Definition 2.3 (Tangent Space $T_{x}X$).
For a smooth manifold $X$ and a point $x\in X$, the tangent space $T_{x}X$ is the vector space of all possible directions one can move from $x$ while remaining on the manifold. Intuitively, if $X$ is a surface in $\mathbb{R}^{3}$, then $T_{x}X$ is the plane tangent to $X$ at $x$. The dimension of $T_{x}X$ equals the dimension of $X$ at $x$.
Definition 2.4 (Transversality).
Let $f:X\to Y$ be a smooth map between smooth manifolds, and let $Z\subset Y$ be a smooth submanifold. We say that $f$ is transverse to $Z$ (written $f\pitchfork Z$) if for every $x\in f^{-1}(Z)$,
$df_{x}(T_{x}X)+T_{f(x)}Z=T_{f(x)}Y,$
where $df_{x}:T_{x}X\to T_{f(x)}Y$ is the derivative of $f$ at $x$. Transversality means that the image of $df_{x}$ and the tangent space of $Z$ together span the entire tangent space of $Y$, ensuring non-degenerate intersections.
Theorem 2.5 (Parametric Transversality Theorem).
Let $F:P\times X\to Y$ be a smooth map, where $P$, $X$, and $Y$ are finite-dimensional smooth manifolds, and let $Z\subset Y$ be a smooth submanifold. If $F$ is transverse to $Z$, then for almost every $p\in P$ (in the sense of Lebesgue measure), the map $F_{p}:X\to Y$ defined by $F_{p}(x)=F(p,x)$ is transverse to $Z$. Moreover, the set $\{p\in P:F_{p}\pitchfork Z\}$ is residual.
Proof.
This is a standard result in differential topology. One proof applies Sard’s theorem to the restriction of the projection $\pi:P\times X\to P$ to the submanifold $F^{-1}(Z)$, which is smooth due to the transversality of $F$. The set of parameters $p$ for which $F_{p}$ is not transverse to $Z$ is contained in the set of critical values of this projection, hence has measure zero. For complete details, see [2, Theorem 8.4] or [13, Chapter 3, Theorem 2.7]. ∎
Corollary 2.6 (Genericity of Transverse Intersections).
Let $\mathcal{F}=\{f_{\mathbf{c}}:X\to Y\}_{\mathbf{c}\in P}$ be a smooth family of maps parameterized by $P=\mathbb{R}^{N}$. If the evaluation map
$F:P\times X\to Y,\quad F(\mathbf{c},x)=f_{\mathbf{c}}(x)$
is transverse to $Z\subset Y$, then for almost every $\mathbf{c}\in\mathbb{R}^{N}$, $f_{\mathbf{c}}$ is transverse to $Z$.
Proof.
This follows directly from Theorem 2.5 since $\mathbb{R}^{N}$ is finite-dimensional and “almost every” refers to Lebesgue measure. ∎
In our context, $X=\mathbb{T}^{d}$ is the torus, $Y=\mathbb{R}^{m}$ for appropriate $m$ (e.g., $m=d+1$ for critical point analysis), and $P=\mathbb{R}^{N}$ is the space of trigonometric polynomial coefficients. The submanifold $Z$ typically represents a set where degeneracies occur (e.g., where critical points become non-Morse). The transversality of $F$ ensures that for generic $\mathbf{c}$, the potential $V(\cdot;\mathbf{c})$ avoids these degeneracies. This is used extensively in Section 3 to prove genericity of Morse properties and unique extrema.
3 Proof of the Main Results
The proof relies on two key lemmas.
Lemma 3.1 (Genericity of Morse Properties and Unique Extrema).
The set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which $V$ is not a Morse function or does not have unique global extrema has Lebesgue measure zero.
Proof.
We prove the lemma using the parametric transversality theory, which are well-suited for analyzing generic properties of families of functions. Let
$V(x;\mathbf{c})=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x}$ (3.1)
be the trigonometric polynomial potential, where $x\in\mathbb{T}^{d}$ and $\mathbf{c}\in\mathbb{R}^{N}$. We treat $V$ as a smooth (in fact, analytic) function in both $x$ and $\mathbf{c}$.
Part 1. Genericity of the Morse Property.
A function $V(\cdot;\mathbf{c})$ is Morse if all its critical points are non-degenerate, i.e., for every $x$ such that $\nabla_{x}V(x;\mathbf{c})=0$, the Hessian matrix $H(V)(x;\mathbf{c})$ is non-singular. To prove that the set
$\mathcal{B}^{(i)}=\{\mathbf{c}\in\mathbb{R}^{N}\mid V(\cdot;\mathbf{c})\text{ is not Morse}\}$
has measure zero, we employ the Parametric Transversality Theorem. Consider the map
$\Phi:\mathbb{T}^{d}\times\mathbb{R}^{N}\to\mathbb{R}^{d}\times\mathbb{R},\quad\Phi(x,\mathbf{c})=\left(\nabla_{x}V(x;\mathbf{c}),\det(H(V)(x;\mathbf{c}))\right).$
A point $(x,\mathbf{c})$ is a degenerate critical point if $\Phi(x,\mathbf{c})=(0,0)$. The set of degenerate critical points is $\mathcal{D}=\Phi^{-1}(0,0)$. By Theorem 2.5, if $\Phi$ is transverse to the submanifold $\{0\}\times\{0\}\subset\mathbb{R}^{d}\times\mathbb{R}$ (i.e., the derivative $D\Phi$ is surjective at points in $\mathcal{D}$), then for almost every $\mathbf{c}\in\mathbb{R}^{N}$, the function $V(\cdot;\mathbf{c})$ has no degenerate critical points, meaning it is Morse.
Since $V$ is analytic in $x$ and linear in $\mathbf{c}$, the map $\Phi$ is smooth. To verify the transversality condition, we explicitly compute the derivative $D\Phi$ with respect to $\mathbf{c}$. For fixed $(x,\mathbf{c})$, the derivative $D_{\mathbf{c}}\Phi(x,\mathbf{c}):\mathbb{R}^{N}\to\mathbb{R}^{d}\times\mathbb{R}$ acts on a vector $v=(v_{m})\in\mathbb{R}^{N}$ as follows:
$D_{\mathbf{c}}\Phi(x,\mathbf{c})[v]$ $=\left(D_{\mathbf{c}}(\nabla_{x}V)(x;\mathbf{c})[v],\ D_{\mathbf{c}}(\det H(V))(x;\mathbf{c})[v]\right)$
$=\left(\nabla_{x}\left(\frac{\partial V}{\partial\mathbf{c}}[v]\right),\ \operatorname{tr}\left(\operatorname{adj}(H(V))\cdot D_{\mathbf{c}}H(V)[v]\right)\right),$ (3.2)
where $\frac{\partial V}{\partial\mathbf{c}}[v]=\sum_{m}v_{m}e^{2\pi im\cdot x}$ is the directional derivative of $V$ with respect to $\mathbf{c}$ in the direction $v$, and we used the identity for the derivative of the determinant.
The key observation is that the components of $\nabla_{x}V$ and $\det H(V)$ are linear in $\mathbf{c}$ with coefficients given by the Fourier basis functions and their derivatives. Specifically, for each coordinate $i=1,\dots,d$:
$\frac{\partial V}{\partial x_{i}}=\sum_{m}c_{m}(2\pi im_{i})e^{2\pi im\cdot x},$
and the Hessian determinant $\det H(V)$ is a homogeneous polynomial of degree $d$ in the second derivatives $\frac{\partial^{2}V}{\partial x_{i}\partial x_{j}}$, which are themselves linear in $\mathbf{c}$ with coefficients involving monomials in $m$.
The linear independence of the Fourier basis $\{e^{2\pi im\cdot x}\}$ over $\mathbb{C}$ (and hence over $\mathbb{R}$ for real parts) implies that the set of functions
$\left\{\frac{\partial V}{\partial x_{i}}(\cdot;\mathbf{c}),\det H(V)(\cdot;\mathbf{c})\right\}$
has full rank for generic $x$ when considered as linear functionals in $\mathbf{c}$. More precisely, the variational derivative of $\Phi$ with respect to $\mathbf{c}$ at a point $(x,\mathbf{c})$ is given by the Jacobian matrix whose columns are the partial derivatives with respect to each $c_{m}$. For the first component:
$\frac{\partial}{\partial c_{m}}\left(\frac{\partial V}{\partial x_{i}}\right)=2\pi im_{i}e^{2\pi im\cdot x},$
and for the second component, by the chain rule:
$\frac{\partial}{\partial c_{m}}(\det H(V))=\operatorname{tr}\left(\operatorname{adj}(H(V))\cdot\frac{\partial H(V)}{\partial c_{m}}\right),$
where $\frac{\partial H(V)}{\partial c_{m}}$ is the Hessian of the basis function $e^{2\pi im\cdot x}$. Since the Fourier basis is linearly independent, the vectors
$\left(2\pi im_{i}e^{2\pi im\cdot x}\right)_{i=1}^{d}\quad\text{and}\quad\operatorname{tr}\left(\operatorname{adj}(H(V))\cdot\frac{\partial H(V)}{\partial c_{m}}\right)$
collectively span $\mathbb{R}^{d}\times\mathbb{R}$ as $m$ varies over the index set, for generic $x$ where $H(V)$ is non-degenerate. This follows because the number of independent Fourier modes exceeds the dimension $d+1$ of the target space.
Thus, $D_{\mathbf{c}}\Phi$ is surjective at points where $\Phi(x,\mathbf{c})=(0,0)$, ensuring transversality. By the parametric transversality theorem, the set of $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ fails to be Morse has measure zero. Hence, $\mathcal{B}^{(i)}$ has measure zero.
Part 2. Genericity of Unique Global Extrema.
We now show that the set
$\mathcal{B}^{(ii)}=\{\mathbf{c}\in\mathbb{R}^{N}\mid V(\cdot;\mathbf{c})\text{ lacks unique global extrema}\}$
has measure zero. A function $V(\cdot;\mathbf{c})$ has unique global extrema if the global maximum and minimum values are attained at unique points in $\mathbb{T}^{d}$. Since $\mathbb{T}^{d}$ is compact, global extrema exist and are critical points. If $V(\cdot;\mathbf{c})$ is Morse (which holds for almost every $\mathbf{c}$ by Part 1), then all critical points are isolated. However, uniqueness of global extrema requires that no two distinct critical points share the same extreme value.
To prove this, we employ a transversality argument on the critical set. Let $\mathcal{C}\subset\mathbb{T}^{d}\times\mathbb{R}^{N}$ be the set of critical points, defined by:
$\mathcal{C}=\{(x,\mathbf{c})\in\mathbb{T}^{d}\times\mathbb{R}^{N}\mid\nabla_{x}V(x;\mathbf{c})=0\}.$
Since $V$ is analytic in $x$ and linear in $\mathbf{c}$, the implicit function theorem implies that $\mathcal{C}$ is a smooth manifold of dimension $N$ (the same as the parameter space) for generic $\mathbf{c}$.
Define the evaluation map on the critical set:
$\Psi:\mathcal{C}\to\mathbb{R},\quad\Psi(x,\mathbf{c})=V(x;\mathbf{c}).$
The global extrema are unique if the restriction of $\Psi$ to each fiber $\mathcal{C}_{\mathbf{c}}=\{x\mid(x,\mathbf{c})\in\mathcal{C}\}$ is injective on the set of extreme values. This is equivalent to requiring that for any two distinct critical points $x\neq y$ in $\mathcal{C}_{\mathbf{c}}$, we have $V(x;\mathbf{c})\neq V(y;\mathbf{c})$.
To analyze this condition, consider the double critical set and the difference map:
$\mathcal{C}^{(2)}=\{((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{C}\times\mathcal{C}\mid x\neq y\},$
which is a smooth manifold of dimension 2Nd2N-d (since the constraint $x\neq y$ removes a submanifold of dimension N1N-1, and $\dim\mathcal{C}=N$). Define the difference map:
$\Xi:\mathcal{C}^{(2)}\to\mathbb{R},\quad\Xi((x,\mathbf{c}),(y,\mathbf{c}))=V(x;\mathbf{c})-V(y;\mathbf{c}).$
The set of parameters $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ has non-unique extrema is exactly the projection onto $\mathbb{R}^{N}$ of the set:
$\mathcal{E}=\Xi^{-1}(0)=\{((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{C}^{(2)}\mid V(x;\mathbf{c})=V(y;\mathbf{c})\}.$
We now show that $\Xi$ is transverse to $\{0\}\subset\mathbb{R}$ (i.e., $D\Xi$ is surjective at points where $\Xi=0$). For a point $((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{E}$, the derivative $D\Xi$ acts on tangent vectors $(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})\in T_{(x,\mathbf{c})}\mathcal{C}\times T_{(y,\mathbf{c})}\mathcal{C}$ as:
$D\Xi[(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})]=\langle\nabla_{x}V(x;\mathbf{c}),\dot{x}\rangle+\langle D_{\mathbf{c}}V(x;\mathbf{c}),\dot{\mathbf{c}}_{1}\rangle-\langle\nabla_{y}V(y;\mathbf{c}),\dot{y}\rangle-\langle D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}_{2}\rangle.$
Since $(x,\mathbf{c}),(y,\mathbf{c})\in\mathcal{C}$, we have $\nabla_{x}V(x;\mathbf{c})=0$ and $\nabla_{y}V(y;\mathbf{c})=0$, so the terms involving $\dot{x}$ and $\dot{y}$ vanish. Thus:
$D\Xi[(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})]=\langle D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}\rangle,$
where $\dot{\mathbf{c}}=\dot{\mathbf{c}}_{1}=\dot{\mathbf{c}}_{2}$ since $\mathbf{c}$ is the same for both points. The linear independence of the Fourier basis $\{e^{2\pi im\cdot x}\}$ ensures that for generic $x\neq y$, the vectors $D_{\mathbf{c}}V(x;\mathbf{c})$ and $D_{\mathbf{c}}V(y;\mathbf{c})$ are linearly independent. Specifically:
$D_{\mathbf{c}}V(x;\mathbf{c})=\left(e^{2\pi im\cdot x}\right)_{|m|\leq n},\quad D_{\mathbf{c}}V(y;\mathbf{c})=\left(e^{2\pi im\cdot y}\right)_{|m|\leq n},$
and since $x\neq y$, the exponential functions are linearly independent for distinct frequencies. Thus, $D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c})\neq 0$ for generic $x\neq y$, ensuring that $D\Xi$ is surjective onto $\mathbb{R}$.
By Theorem 2.5, the set $\mathcal{E}$ is a smooth submanifold of $\mathcal{C}^{(2)}$ with codimension 1. The projection $\pi:\mathcal{C}^{(2)}\to\mathbb{R}^{N}$ given by $\pi((x,\mathbf{c}),(y,\mathbf{c}))=\mathbf{c}$ is a smooth map. Since $\dim\mathcal{E}=\dim\mathcal{C}^{(2)}-1=2N-d-1$ and $\dim\mathbb{R}^{N}=N$, if $N>d+1$ (which holds for sufficiently high-degree trigonometric polynomials), then $\dim\mathcal{E}<N$, and by Sard’s theorem, the set of critical values of $\pi|_{\mathcal{E}}$ has measure zero in $\mathbb{R}^{N}$. The set of $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ has non-unique extrema is contained in this set of critical values, hence has measure zero. Therefore, $\mathcal{B}^{(ii)}$ has measure zero.
Conclusion.
Having shown that both $\mathcal{B}^{(i)}$ (non-Morse potentials) and $\mathcal{B}^{(ii)}$ (potentials without unique extrema) are subsets of parameter sets where the respective transversality conditions fail (Theorems 2.5), and since such failure sets are smooth submanifolds of positive codimension (hence of Lebesgue measure zero), their union $\mathcal{B}^{(i)}\cup\mathcal{B}^{(ii)}$ also has measure zero. This completes the proof that for almost every $\mathbf{c}\in\mathbb{R}^{N}$, the potential $V(\cdot;\mathbf{c})$ is a Morse function with unique global extrema. ∎
Lemma 3.2 (Genericity of Cartan Estimates).
The set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which the potential $V(\cdot;\mathbf{c})$ fails to satisfy the Cartan estimates (conditions (iii) and (iv) of Definition 2.1) has Lebesgue measure zero.
Proof.
We prove the lemma using Cartan-type estimates for analytic functions and a parameter-space argument. Let
$V(x;\mathbf{c})=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x}$
be the trigonometric polynomial potential, where $\mathbf{c}=(c_{m})\in\mathbb{R}^{N}$. For fixed $h\in\mathbb{T}^{d}$, define the analytic functions:
$F_{h}(x;\mathbf{c})=V(x+h;\mathbf{c})-V(x;\mathbf{c}),\quad G_{h}(x;\mathbf{c})=g_{V,h,i,j}(x;\mathbf{c}),$ (3.3)
where $g_{V,h,i,j}$ is the determinant function derived from $V$ as in condition (iii) of class $\mathfrak{G}$. The Cartan estimate requires that for any $K\gg 1$ and $\|h\|\geq\exp(-K)$,
$\text{mes}\left\{x\in\mathbb{T}^{d}:\min\left(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|\right)<\exp(-K)\right\}\leq\exp(-K^{\mathfrak{c}_{1}}),$ (3.4)
where $\mathfrak{c}_{1}>0$ is a constant. Define the bad set for fixed $h$ and $K$ as:
$\mathcal{B}^{(iii)}(h,K)=\left\{\mathbf{c}\in\mathbb{R}^{N}:\text{mes}\left\{x:\min\left(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}})\right\}.$ (3.5)
We aim to show that the overall bad set $\mathcal{B}^{(iii)}=\bigcup_{h\in\mathbb{T}^{d}}\bigcup_{K\geq K_{0}}\mathcal{B}^{(iii)}(h,K)$ has measure zero.
The proof proceeds in three steps:
(1) Uniform Cartan estimate for bounded parameters: For any fixed $R>0$, consider the ball $B_{R}=\{\mathbf{c}\in\mathbb{R}^{N}:\|\mathbf{c}\|\leq R\}$. Since $F_{h}$ and $G_{h}$ are analytic in $x$ and linear in $\mathbf{c}$, the family $\{F_{h}(\cdot;\mathbf{c}),G_{h}(\cdot;\mathbf{c}):\mathbf{c}\in B_{R}\}$ is uniformly bounded and equicontinuous on $\mathbb{T}^{d}$. By the Cartan-type estimate for analytic functions (see Cartan [7] or Goldstein et al. [11]), there exist constants $C_{R}>0$ and $\delta_{R}>0$ (depending on $R$ and $h$) such that for all $\mathbf{c}\in B_{R}$ with $F_{h}$ and $G_{h}$ not both identically zero,
$\text{mes}\left\{x:\min\left(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|\right)<\epsilon\right\}\leq C_{R}\epsilon^{\delta_{R}}.$ (3.6)
To derive this, recall that for an analytic function $f$ on $\mathbb{T}^{d}$ with analytic norm $\|f\|_{\mathcal{A}}$, the Cartan estimate guarantees that for any $\epsilon>0$,
$\text{mes}\left\{x:|f(x)|<\epsilon\right\}\leq C\exp\left(-\frac{\delta}{\|f\|_{\mathcal{A}}}\log\frac{1}{\epsilon}\right),$ (3.7)
where $C,\delta>0$ depend on the dimension $d$. Applying this to $f=\min(|F_{h}|,|G_{h}|)$ and noting that $\|F_{h}\|_{\mathcal{A}}\leq 2\|V\|_{\mathcal{A}}$ and $\|G_{h}\|_{\mathcal{A}}\leq C^{\prime}\|V\|_{\mathcal{A}}$ for some $C^{\prime}>0$, we obtain (3.6) with $\delta_{R}$ proportional to $1/\|V\|_{\mathcal{A}}$. Since $\|V\|_{\mathcal{A}}$ is uniformly bounded for $\mathbf{c}\in B_{R}$, the constants $C_{R}$ and $\delta_{R}$ are uniform over $B_{R}$.
(2) Measure zero for bounded parameters:. Now, fix $h$ and $R$. The set of $\mathbf{c}\in B_{R}$ for which $F_{h}$ or $G_{h}$ is identically zero is contained in a proper linear subspace of $\mathbb{R}^{N}$. Specifically, if $F_{h}\equiv 0$, then for all $x$,
$\sum_{m}c_{m}\left(e^{2\pi im\cdot(x+h)}-e^{2\pi im\cdot x}\right)=0,$ (3.8)
which implies that $c_{m}(e^{2\pi im\cdot h}-1)=0$ for all $m$. For generic $h$, $e^{2\pi im\cdot h}\neq 1$ for $m\neq 0$, so $c_{m}=0$ for $m\neq 0$. Thus, the set of such $\mathbf{c}$ has measure zero. Similarly for $G_{h}$. Hence, for $\mathbf{c}\in B_{R}$ outside this null set, the Cartan estimate (3.6) applies. Choose $K_{0}(R,h)$ large enough such that for all $K\geq K_{0}(R,h)$, we have
$C_{R}\exp(-\delta_{R}K)\leq\exp(-K^{\mathfrak{c}_{1}}).$ (3.9)
Then for $K\geq K_{0}(R,h)$, the bad set $\mathcal{B}^{(iii)}(h,K)\cap B_{R}$ is empty. For $K<K_{0}(R,h)$, the bad set is contained in the null set where $F_{h}$ or $G_{h}$ is identically zero. Thus, for fixed $h$ and $R$, the set $\bigcup_{K\geq K_{0}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}$ has measure zero.
(3) Union over $h$ and $K$, and unbounded parameters:. To handle the union over $h\in\mathbb{T}^{d}$, we use a covering argument. For each $K\geq K_{0}$, cover $\mathbb{T}^{d}$ with a grid of points $h_{j}$ such that the number of points is $N(K)=O(\exp(Kd))$ and the mesh size is $\exp(-K)$. By the uniform continuity of $F_{h}$ and $G_{h}$ in $h$, for any $h$ in the $\exp(-K)$-neighborhood of $h_{j}$, we have
$|F_{h}(x;\mathbf{c})-F_{h_{j}}(x;\mathbf{c})|\leq L\exp(-K),\quad|G_{h}(x;\mathbf{c})-G_{h_{j}}(x;\mathbf{c})|\leq L\exp(-K),$ (3.10)
where $L$ is a Lipschitz constant depending on $R$. Thus, if
$\min(|F_{h_{j}}(x;\mathbf{c})|,|G_{h_{j}}(x;\mathbf{c})|)\geq 2L\exp(-K),$
then $\min(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|)\geq L\exp(-K)$. This allows us to transfer the Cartan estimate from $h_{j}$ to all $h$ in its neighborhood. Specifically, for each $h_{j}$, we have from Step (2) that
$\text{mes}\left(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R}\right)=0\quad\text{for }K\geq K_{0}(R,h_{j}).$ (3.11)
By the covering, the union over $h$ can be controlled by the sum over $j$:
$\text{mes}\left(\bigcup_{h\in\mathbb{T}^{d}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}\right)\leq\sum_{j=1}^{N(K)}\text{mes}\left(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R}\right).$ (3.12)
From (3.6) and the choice of $K_{0}$, for $K\geq\max_{j}K_{0}(R,h_{j})$, each term in the sum is zero. For $K<\max_{j}K_{0}(R,h_{j})$, we use a coarse bound: since $\mathcal{B}^{(iii)}(h_{j},K)\subset B_{R}$ and $\text{mes}(B_{R})<\infty$, we have
$\text{mes}\left(\bigcup_{h\in\mathbb{T}^{d}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}\right)\leq N(K)\cdot\text{mes}(B_{R})\leq M\exp(Kd)\cdot\exp(-C^{\prime}K)$ (3.13)
for some $M,C^{\prime}>0$, where $C^{\prime}$ is chosen such that $\exp(-C^{\prime}K)$ dominates the measure of the bad set for each $h_{j}$. Since
$\operatorname{mes}(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R})\leq\exp(-C^{\prime}K)$
for some $C^{\prime}>d$ (from the Cartan estimate), the series
$\sum_{K\geq K_{0}}\operatorname{mes}(\mathcal{B}^{(iii)}(h_{j},K)\cap B_{R})$
converges. By the Borel-Cantelli lemma, the set
$\bigcup_{K\geq K_{0}}\bigcup_{h\in\mathbb{T}^{d}}\mathcal{B}^{(iii)}(h,K)\cap B_{R}$
has measure zero. Since $\mathbb{R}^{N}=\bigcup_{R\in\mathbb{N}}B_{R}$ and measure zero is preserved under countable unions, the full bad set $\mathcal{B}^{(iii)}$ has measure zero.
We now turn to the genericity of the gradient Cartan estimate (condition (iv)). Let $\mathcal{B}^{(iv)}$ be the set of all coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which the gradient Cartan estimate fails, i.e., there exists some $\eta\in\mathbb{R}$, some unit vector $h_{0}\in\mathbb{R}^{d}$, and some $K\geq K_{0}$ such that:
$\text{mes}\left\{x\in\mathbb{T}^{d}:\min\left(|V(x;\mathbf{c})-\eta|,|\langle\nabla V(x;\mathbf{c}),h_{0}\rangle|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}}).$
More formally, we can express this as:
$\mathcal{B}^{(iv)}=\bigcup_{h_{0}\in S^{d-1}}\bigcup_{\eta\in\mathbb{R}}\bigcup_{K\geq K_{0}}\mathcal{B}^{(iv)}(\eta,h_{0},K)$
where $S^{d-1}$ is the unit sphere in $\mathbb{R}^{d}$, and for fixed $\eta$, $h_{0}$, and $K$, the set $\mathcal{B}^{(iv)}(\eta,h_{0},K)$ is defined as:
$\mathcal{B}^{(iv)}(\eta,h_{0},K)$ (3.14)
$=\left\{\mathbf{c}\in\mathbb{R}^{N}:\text{mes}\left\{x:\min\left(|V(x;\mathbf{c})-\eta|,|\langle\nabla V(x;\mathbf{c}),h_{0}\rangle|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}})\right\}.$ (3.15)
The argument for showing that $\mathcal{B}^{(iv)}$ has measure zero parallels that for $\mathcal{B}^{(iii)}$ but requires careful treatment of the additional parameters $\eta$ and $h_{0}$.
For the analysis, it is convenient to define the relevant analytic functions:
$F_{\eta}(x;\mathbf{c})=V(x;\mathbf{c})-\eta,\quad G_{h_{0}}(x;\mathbf{c})=\langle\nabla V(x;\mathbf{c}),h_{0}\rangle.$ (3.16)
In terms of these functions, the gradient Cartan estimate requires that for any $K\gg 1$,
$\text{mes}\left\{x\in\mathbb{T}^{d}:\min\left(|F_{\eta}(x;\mathbf{c})|,|G_{h_{0}}(x;\mathbf{c})|\right)<\exp(-K)\right\}\leq\exp(-K^{\mathfrak{c}_{1}}).$ (3.17)
The corresponding bad set for fixed $\eta$, $h_{0}$, and $K$ is:
$\mathcal{B}^{(iv)}(\eta,h_{0},K)$ (3.18)
$=\left\{\mathbf{c}\in\mathbb{R}^{N}:\text{mes}\left\{x:\min\left(|F_{\eta}(x;\mathbf{c})|,|G_{h_{0}}(x;\mathbf{c})|\right)<\exp(-K)\right\}>\exp(-K^{\mathfrak{c}_{1}})\right\}.$ (3.19)
The overall bad set is
$\mathcal{B}^{(iv)}=\bigcup_{h_{0}\in S^{d-1}}\bigcup_{\eta\in\mathbb{R}}\bigcup_{K\geq K_{0}}\mathcal{B}^{(iv)}(\eta,h_{0},K),$
where $S^{d-1}$ is the unit sphere.
Since $\sum_{K\geq K_{0}}\exp(-(C^{\prime}-d)K)<\infty$ when $C^{\prime}>d$, by the Borel–Cantelli lemma, the union of the bad sets has measure zero.
The proof proceeds using the same three-step structure as for $\mathcal{B}^{(iii)}$:
(4) Uniform Estimate for Bounded Parameters: For a fixed radius $R>0$ and parameters $\mathbf{c}$ in the ball $B_{R}=\{\mathbf{c}\in\mathbb{R}^{N}:\|\mathbf{c}\|\leq R\}$, the functions $F_{\eta}$ and $G_{h_{0}}$ are analytic in $x$ and linear in $\mathbf{c}$. Applying the Cartan-type estimate yields constants $\tilde{C}_{R},\tilde{\delta}_{R}>0$ such that for $\mathbf{c}\in B_{R}$ (excluding a set where $F_{\eta}$ and $G_{h_{0}}$ are identically zero, which has measure zero),
$\text{mes}\left\{x:\min\left(|F_{\eta}(x;\mathbf{c})|,|G_{h_{0}}(x;\mathbf{c})|\right)<\epsilon\right\}\leq\tilde{C}_{R}\epsilon^{\tilde{\delta}_{R}}.$ (3.20)
(5) Measure Zero for Bounded Parameters: For fixed $\eta$, $h_{0}$, and $R$, we choose $K_{0}(R,\eta,h_{0})$ large enough so that for $K\geq K_{0}$,
$\tilde{C}_{R}\exp(-\tilde{\delta}_{R}K)\leq\exp(-K^{\mathfrak{c}_{1}}).$
Then, for such $K$,
$\mathcal{B}^{(iv)}(\eta,h_{0},K)\cap B_{R}=\emptyset.$
For $K<K_{0}$, the bad set is contained in the null set where the functions are identically zero. Thus, $\bigcup_{K\geq K_{0}}\mathcal{B}^{(iv)}(\eta,h_{0},K)\cap B_{R}$ has measure zero.
(6) Union over $\eta$, $h_{0}$, $K$, and Unbounded Parameters: This is the most critical step, addressing the union over the non-compact set $\mathbb{R}$ for $\eta$ and the compact set $S^{d-1}$ for $h_{0}$.
  • Union over $\eta$: Although $\eta\in\mathbb{R}$ is non-compact, the function $V(x;\mathbf{c})$ is bounded for $\mathbf{c}\in B_{R}$. Let $I_{R}$ be a compact interval containing the range of $V$ for all $\mathbf{c}\in B_{R}$, e.g.,
    $I_{R}=[-\max_{\mathbf{c}\in B_{R}}\|V(\cdot;\mathbf{c})\|_{\infty},\max_{\textbf{c}\in B_{R}}\|V(\cdot;\textbf{ c})\|_{\infty}].$
    If $\eta\notin I_{R}$, then $|V(x;\mathbf{c})-\eta|$ is bounded away from zero uniformly, and the Cartan estimate holds trivially. Therefore, it suffices to consider $\eta\in I_{R}$, a compact set. We can cover $I_{R}$ with $O(\exp(K))$ intervals of length $\exp(-K)$.
  • Union over $h_{0}$: The unit sphere $S^{d-1}$ is compact. We cover it with a grid of $O(\exp(K(d-1)))$ points $h_{0,j}$ with mesh size $\exp(-K)$.
  • Using the uniform continuity of the functions with respect to $\eta$ and $h_{0}$, the estimate for a fixed $(\eta_{j},h_{0,k})$ can be extended to a neighborhood of size $\exp(-K)$ around them, with a possible adjustment of constants.
Following the same covering argument and application of the Borel-Cantelli lemma as in Step (3) of the original proof, and then taking the union over $R\in\mathbb{N}$, we conclude that the full set $\mathcal{B}^{(iv)}$ has Lebesgue measure zero.
Therefore, the set of coefficients for which any Cartan estimate fails has Lebesgue measure zero. ∎
Proof of Theorem 2.2.
The theorem asserts that the set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which $V(\cdot;\mathbf{c})\in\mathfrak{G}$ has full Lebesgue measure. Recall that $\mathfrak{G}$ is defined by four conditions (Definition 2.1). Let $\mathcal{B}=\mathcal{B}^{(i)}\cup\mathcal{B}^{(ii)}\cup\mathcal{B}^{(iii)}\cup\mathcal{B}^{(iv)}$ be the set of parameters where at least one condition fails.
By Lemma 3.1, the sets $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ have measure zero. By Lemma 3.2, the sets $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ also have measure zero. Since a finite union of measure zero sets has measure zero, $\mathcal{B}$ has measure zero. Consequently, the set of good parameters $\mathcal{G}=\mathbb{R}^{N}\setminus\mathcal{B}$ has full Lebesgue measure. This completes the proof of the genericity conjecture by Goldstein et al. [11]. ∎
Remark 3.3 (Topological Genericity).
The proof of Theorem 2.2 actually establishes a stronger result: the set of good parameters $\mathcal{G}$ is not only of full measure but also a residual set in $\mathbb{R}^{N}$. This follows because the exceptional sets $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$, as shown in Lemma 3.1, are contained in smooth submanifolds of positive codimension, which are closed and have empty interior (hence nowhere dense). The exceptional sets $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ from Lemma 3.2 can also be shown to be of first category. Thus, $\mathcal{G}$ is the complement of a finite union of first category sets, making it a residual set. This confirms that the potentials in class $\mathfrak{G}$ are generic in both the measure-theoretic and topological senses.
Acknowledgments
This work was supported by NSFC (No. 11571327, 11971059)