### v2

# On the Genericity of the Spectrum Intervalization for Multi-Frequency Quasiperiodic Schrödinger Operators

Abstract
This paper proves a genericity conjecture by Goldstein, Schlag, and Voda [Invent. Math. 217 (2019)] for multi-frequency quasiperiodic Schrödinger operators. Specifically, we show that for almost all coefficients of real trigonometric polynomial potentials, the spectrum forms a single interval under strong coupling conditions. This confirms a long-standing intuition by Chulaevsky and Sinai [Comm. Math. Phys. 125 (1989)] that the spectrum typically consists of an interval for generic potentials, and extends the existence results of Goldstein et al. to a full measure setting. Our proof relies on tools from differential topology, measure theory, and analytic function theory.
Mathematics Subject Classification (2020): 47A10, 47B39
keywords:
Quasiperiodic Schrödinger operators , Spectrum , Genericity , Cartan estimates , Transversality theory.
journal: arXiv
1 Introduction
Quasiperiodic Schrödinger operators have been extensively studied in mathematical physics, particularly in the context of Anderson localization [5, 4, 7] and spectral theory [6, 10]. A fundamental question concerns the structure of the spectrum: whether it is a Cantor set [1] or a single interval [13]. For multi-frequency operators with analytic potentials, it was first suggested by Chulaevsky and Sinai [9] that under strong coupling, the spectrum typically forms an interval for generic potentials. This intuition was later formalized by Goldstein, Schlag, and Voda [13], who proved that for a specific class of potentials (denoted class $\mathfrak{G}$), the spectrum is indeed an interval. Moreover, they conjectured that class $\mathfrak{G}$ is generic, i.e., holds for almost all coefficients in the space of trigonometric polynomials.
In this paper, we prove the genericity conjecture by Goldstein et al. [13], showing that for almost all coefficients of real trigonometric polynomials, the potential belongs to class $\mathfrak{G}$, which ensures the spectrum is a single interval. Our approach combines tools from differential topology, measure theory, and analytic function theory, specifically leveraging the parametric transversality theorem and Cartan-type estimates to establish the full-measure property of class $\mathfrak{G}$. This provides a comprehensive framework for understanding the generic behavior of these operators and confirms the original intuition of Chulaevsky and Sinai [9].
2 Preliminaries
2.1 Class $\mathfrak{G}$ of Potentials
Consider the multi-frequency quasiperiodic Schrödinger operator on $\ell^{2}(\mathbb{Z})$:
$(H(x)\psi)(n)=-\psi(n+1)-\psi(n-1)+\lambda V(x+n\omega)\psi(n),$ (2.1)
where $x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}$ is a Diophantine frequency vector, $\lambda>0$ is the coupling constant, and $V$ is a real analytic potential.
We focus on trigonometric polynomial potentials of the form:
$V(x)=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x},$ (2.2)
with coefficients $\mathbf{c}=(c_{m})\in\mathbb{R}^{N}$, where NN is the number of integer vectors $m\in\mathbb{Z}^{d}$ satisfying $|m|=\sum_{j=1}^{d}|m_{j}|\leq n$. A combinatorial count gives
$N=\sum_{k=0}^{n}a_{d,k},\quad\text{where}\quad a_{d,k}=\sum_{l=1}^{\min(d,k)}\binom{d}{l}2^{l}\binom{k-1}{l-1}\ \text{for}\ k\geq 1,\ \text{and}\ a_{d,0}=1.$ (2.3)
Goldstein et al. [13] introduced the following function class:
Definition 2.1 (Class $\mathfrak{G}$).
A potential $V$ belongs to class $\mathfrak{G}$ if it satisfies the following conditions:
  1. (i)
    $V$ is a Morse function (all critical points are non-degenerate).
  2. (ii)
    $V$ has unique global minimum and maximum points.
  3. (iii)
    For any $h\in\mathbb{T}^{d}$ with $\|h\|\geq\exp(-K)$ and $K\gg 1$, the Cartan estimate holds:
    ${\rm mes}\left\{x\in\mathbb{T}^{d}:\min\left(|V(x+h)-V(x)|,|g_{V,h,i,j}(x)|\right)<\exp(-K)\right\}$
    $\quad\leq\exp(-K^{\mathfrak{c}_{1}}),$ (2.4)
    where
    $g_{V,h,i,j}(x)=\operatorname{det}\begin{bmatrix}\partial_{x_{i}}V(x)&\partial_{x_{j}}V(x)\\ \partial_{x_{i}}V(x+h)&\partial_{x_{j}}V(x+h)\end{bmatrix}$.
    (see [8] for the origin of such estimates). The constant $\mathfrak{c}_{1}>0$ is the same as in [13].
  4. (iv)
    For any unit vector $h_{0}\in\mathbb{R}^{d}$ and $\eta\in\mathbb{R}$, the gradient Cartan estimate holds:
    ${\rm mes}\left\{x\in\mathbb{T}^{d}:\min\left(|V(x)-\eta|,|\langle\nabla V(x),h_{0}\rangle|\right)<\exp(-K)\right\}\leq\exp(-K^{\mathfrak{c}_{1}}).$ (2.5)
2.2 Statement of Main Theorem
We now state the main theorem, which proves the genericity conjecture by Goldstein et al. [13, Remark 1.2(b)].
Theorem 2.2 (Genericity of Class $\mathfrak{G}$).
Let $V$ be a trigonometric polynomial potential as in (2.2). Then the set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which $V$ belongs to class $\mathfrak{G}$ has full Lebesgue measure. Consequently, for almost all $\mathbf{c}$, under strong coupling ($\lambda\gg 1$), the spectrum of the operator $H(x)$ is a single interval.
The proof of Theorem 2.2 relies on two key lemmas: Lemma 3.1, which uses the parametric transversality theorem to establish genericity of Morse functions and unique extrema, and Lemma 3.6, which uses Cartan estimates and Borel-Cantelli arguments to show that the required analytic estimates hold for almost every coefficient.
2.3 Parametric Transversality Theorem
We now recall the essential tools from differential topology [2, 14, 15, 16] which are used to prove that key properties hold for almost every potential in our genericity analysis. These results are foundational for establishing that exceptional sets have measure zero in finite-dimensional parameter spaces.
Definition 2.3 (Tangent Space $T_{x}X$).
For a smooth manifold $X$ and a point $x\in X$, the tangent space $T_{x}X$ is the vector space of all possible directions one can move from $x$ while remaining on the manifold. The dimension of $T_{x}X$ equals the dimension of $X$ at $x$.
Definition 2.4 (Transversality).
Let $f:X\to Y$ be a smooth map between smooth manifolds, and let $Z\subset Y$ be a smooth submanifold. We say that $f$ is transverse to $Z$ (written $f\pitchfork Z$) if for every $x\in f^{-1}(Z)$,
$df_{x}(T_{x}X)+T_{f(x)}Z=T_{f(x)}Y,$
where $df_{x}:T_{x}X\to T_{f(x)}Y$ is the derivative of $f$ at $x$.
Theorem 2.5 (Parametric Transversality Theorem).
Let $F:P\times X\to Y$ be a smooth map, where $P$, $X$, and $Y$ are finite-dimensional smooth manifolds, and let $Z\subset Y$ be a smooth submanifold. If $F$ is transverse to $Z$, then for almost every $p\in P$ (in the sense of Lebesgue measure), the map $F_{p}:X\to Y$ defined by $F_{p}(x)=F(p,x)$ is transverse to $Z$. Moreover, the set $\{p\in P:F_{p}\pitchfork Z\}$ is residual.
Proof.
This is a standard result in differential topology. One proof applies Sard’s theorem to the restriction of the projection $\pi:P\times X\to P$ to the submanifold $F^{-1}(Z)$, which is smooth due to the transversality of $F$. The set of parameters $p$ for which $F_{p}$ is not transverse to $Z$ is contained in the set of critical values of this projection, hence has measure zero. For complete details, see [2, Theorem 8.4] or [15, Chapter 3, Theorem 2.7]. ∎
Corollary 2.6 (Genericity of Transverse Intersections).
Let $\mathcal{F}=\{f_{\mathbf{c}}:X\to Y\}_{\mathbf{c}\in P}$ be a smooth family of maps parameterized by $P=\mathbb{R}^{N}$. If the evaluation map
$F:P\times X\to Y,\quad F(\mathbf{c},x)=f_{\mathbf{c}}(x)$
is transverse to $Z\subset Y$, then for almost every $\mathbf{c}\in\mathbb{R}^{N}$, $f_{\mathbf{c}}$ is transverse to $Z$.
Proof.
This follows directly from Theorem 2.5 since $\mathbb{R}^{N}$ is finite-dimensional and “almost every” refers to Lebesgue measure. ∎
In our context, $X=\mathbb{T}^{d}$ is the torus, $Y=\mathbb{R}^{m}$ for appropriate $m$ (e.g., $m=d+1$ for critical point analysis), and $P=\mathbb{R}^{N}$ is the space of trigonometric polynomial coefficients. The submanifold $Z$ typically represents a set where degeneracies occur (e.g., where critical points become non-Morse). The transversality of $F$ ensures that for generic $\mathbf{c}$, the potential $V(\cdot;\mathbf{c})$ avoids these degeneracies.
2.4 Borel-Cantelli Lemma
Proposition 2.7 (Borel-Cantelli Lemma).
Let $(\Omega,\mathcal{F},\mu)$ be a measure space with $\mu(\Omega)<\infty$ (or more generally a probability space).
  1. (1)
    First Borel-Cantelli Lemma: If $\{A_{k}\}_{k\geq 1}$ is a sequence of measurable sets such that $\sum_{k=1}^{\infty}\mu(A_{k})<\infty$, then
    $\mu\Bigl(\bigcap_{n=1}^{\infty}\bigcup_{k=n}^{\infty}A_{k}\Bigr)=0.$
    In words, almost every point belongs to only finitely many $A_{k}$.
  2. (2)
    Second Borel-Cantelli Lemma: If $\{A_{k}\}_{k\geq 1}$ are independent events in a probability space and $\sum_{k=1}^{\infty}\mu(A_{k})=\infty$, then
    $\mu\Bigl(\bigcap_{n=1}^{\infty}\bigcup_{k=n}^{\infty}A_{k}\Bigr)=1,$
    i.e., almost every point belongs to infinitely many $A_{k}$.
For a proof see e.g. [11, Theorem 2.3.1 and 2.3.7] or [3, Theorem 4.2 and 4.3].
3 Proof of the Main Results
3.1 Genericity of Morse Properties and Unique Extrema
Lemma 3.1 (Genericity of Morse Properties and Unique Extrema).
The set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which $V$ is not a Morse function or does not have unique global extrema has Lebesgue measure zero.
Proof.
We prove the lemma using algebraic and differential topological methods. Let
$V(x;\mathbf{c})=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x}$
be the trigonometric polynomial potential, where $x\in\mathbb{T}^{d}$ and $\mathbf{c}\in\mathbb{R}^{N}$. $V$ is smooth (in fact, analytic) in both $x$ and $\mathbf{c}$.
Part 1. Genericity of the Morse Property.
A function $V(\cdot;\mathbf{c})$ is Morse iff it has no degenerate critical points, i.e., there is no pair $(x,\mathbf{c})$ with $\nabla_{x}V(x;\mathbf{c})=0$ and $\det H(V)(x;\mathbf{c})=0$. Define
$\Phi:\mathbb{T}^{d}\times\mathbb{R}^{N}\to\mathbb{R}^{d}\times\mathbb{R},\qquad\Phi(x,\mathbf{c})=\bigl(\nabla_{x}V(x;\mathbf{c}),\;\det H(V)(x;\mathbf{c})\bigr).$
The set of degenerate critical points is $\mathcal{D}=\Phi^{-1}(0,0)$. The projection $\pi(x,\mathbf{c})=\mathbf{c}$ sends $\mathcal{D}$ onto $\mathcal{B}^{(i)}$, the set of parameters for which $V(\cdot;\mathbf{c})$ has at least one degenerate critical point. Thus it suffices to prove that $\pi(\mathcal{D})$ has Lebesgue measure zero.
We use an algebraic argument. Because $V$ is a trigonometric polynomial, the equations $\nabla_{x}V(x;\mathbf{c})=0$ are linear in $\mathbf{c}$ with coefficients that are analytic functions of $x$ (specifically, Fourier exponentials). The equation $\det H(V)(x;\mathbf{c})=0$ is a polynomial of degree $d$ in the second derivatives, hence also polynomial in $\mathbf{c}$ with analytic coefficients.
We can eliminate $x$ by expanding everything in the Fourier basis. Write
$\nabla_{x}V(x;\mathbf{c})=\sum_{m}c_{m}(2\pi im)e^{2\pi im\cdot x},$
and similarly for $\det H(V)(x;\mathbf{c})$ which becomes a sum $\sum_{m}c_{m}^{(d)}e^{2\pi im\cdot x}$ where $c_{m}^{(d)}$ are homogeneous polynomials of degree $d$ in the coefficients $\{c_{m}\}$ (coming from products of derivatives). The condition that there exists $x\in\mathbb{T}^{d}$ such that all these Fourier series vanish simultaneously is equivalent to the vanishing of all their Fourier coefficients (since the exponentials are linearly independent). More concretely, let $\{e_{k}(x)\}_{k\in\mathcal{K}}$ be a finite basis of trigonometric polynomials of degree $\leq n$ (e.g., $e^{2\pi im\cdot x}$ with $|m|\leq n$). Then $\nabla_{x}V$ and $\det H(V)$ are finite linear combinations of these basis functions. The existence of an $x$ making them zero is equivalent to the existence of $x$ such that a finite system of analytic equations holds. Eliminating $x$ by using resultants (or by noting that the condition that a finite set of analytic functions has a common zero is itself an analytic condition on the coefficients $\mathbf{c}$) yields a finite set of analytic functions $f_{1}(\mathbf{c}),\dots,f_{r}(\mathbf{c})$ such that
$\pi(\mathcal{D})=\{\mathbf{c}\in\mathbb{R}^{N}:f_{1}(\mathbf{c})=\cdots=f_{r}(\mathbf{c})=0\}.$
In other words, $\pi(\mathcal{D})$ is an analytic subset of $\mathbb{R}^{N}$ (in fact, it is a real algebraic set because all equations are polynomial in $\mathbf{c}$ after clearing denominators, but analytic suffices).
We claim that $\pi(\mathcal{D})$ is not the whole space $\mathbb{R}^{N}$. Indeed, take the specific potential $V_{0}(x)=\sum_{j=1}^{d}\cos(2\pi x_{j})$. Its coefficient vector $\mathbf{c}_{0}$ corresponds to a Morse function (the critical points are at $x_{j}=0$ or $1/2$, and the Hessian is non-degenerate). For this $\mathbf{c}_{0}$, there is no $x$ with $\nabla V=0$ and $\det H(V)=0$ simultaneously. Hence $\mathbf{c}_{0}\notin\pi(\mathcal{D})$, so $\pi(\mathcal{D})\neq\mathbb{R}^{N}$.
A proper analytic subset of $\mathbb{R}^{N}$ has Lebesgue measure zero (it is contained in the zero set of a non-zero analytic function, and the zero set of a non-zero analytic function has measure zero). Therefore $\mathcal{B}^{(i)}=\pi(\mathcal{D})$ has measure zero.
(For readers who prefer an even more elementary argument: For each fixed $x$, the conditions $\nabla_{x}V(x;\mathbf{c})=0$ and $\det H(V)(x;\mathbf{c})=0$ define a proper algebraic subset $A_{x}\subset\mathbb{R}^{N}$ (since they are non-trivial linear/polynomial equations). The union $\bigcup_{x\in\mathbb{T}^{d}}A_{x}$ is the set of $\mathbf{c}$ for which there exists some $x$ with degenerate critical point. However, this union is uncountable and could be large; but one can prove that $\bigcup_{x}A_{x}$ is actually a finite union because the equations are analytic and the set of $x$ that can give a solution for some $\mathbf{c}$ is compact and the condition is semi-algebraic. Alternatively, the earlier elimination argument is standard and rigorous.)
Part 2. Genericity of Unique Global Extrema.
Let
$\mathcal{B}^{(ii)}=\{\mathbf{c}\in\mathbb{R}^{N}\mid V(\cdot;\mathbf{c})\text{ lacks unique global extrema}\}.$
We show that $\mathcal{B}^{(ii)}$ has measure zero. By Part 1, for almost every $\mathbf{c}$, $V(\cdot;\mathbf{c})$ is Morse, hence its critical points are isolated. Global extrema are attained at critical points. Uniqueness fails if there exist two distinct critical points $x\neq y$ with $V(x;\mathbf{c})=V(y;\mathbf{c})$.
Define the critical set
$\mathcal{C}=\{(x,\mathbf{c})\in\mathbb{T}^{d}\times\mathbb{R}^{N}\mid\nabla_{x}V(x;\mathbf{c})=0\}.$
For a fixed $\mathbf{c}$ where $V(\cdot;\mathbf{c})$ is Morse, the set $\mathcal{C}_{\mathbf{c}}=\{x:\nabla_{x}V(x;\mathbf{c})=0\}$ is finite and its cardinality is bounded by a constant depending only on the degree $n$. Moreover, the map $\pi:\mathcal{C}\to\mathbb{R}^{N}$ given by $\pi(x,\mathbf{c})=\mathbf{c}$ has the property that for almost every $\mathbf{c}$ the fiber $\pi^{-1}(\mathbf{c})$ is discrete. By the implicit function theorem, $\mathcal{C}$ is a smooth submanifold of $\mathbb{T}^{d}\times\mathbb{R}^{N}$ of dimension NN (the dimension of the parameter space). The exceptional set where this fails is contained in the set of $\mathbf{c}$ where $V(\cdot;\mathbf{c})$ has a degenerate critical point, which has measure zero by Part 1.
Consider the double critical set
$\mathcal{C}^{(2)}=\{((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{C}\times\mathcal{C}\mid x\neq y\},$
which is a smooth manifold of dimension 2Nd2N-d (the constraint $x\neq y$ removes a submanifold of codimension $d$). Define the difference map
$\Xi:\mathcal{C}^{(2)}\to\mathbb{R},\quad\Xi((x,\mathbf{c}),(y,\mathbf{c}))=V(x;\mathbf{c})-V(y;\mathbf{c}).$
The set of parameters for which $V(\cdot;\mathbf{c})$ has non-unique extrema is contained in the projection $\pi(\mathcal{E})$, where $\mathcal{E}=\Xi^{-1}(0)$ and $\pi((x,\mathbf{c}),(y,\mathbf{c}))=\mathbf{c}$.
We claim that $\Xi$ is transverse to $\{0\}$. Indeed, for $((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{E}$, a tangent vector $(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})$ in $T_{(x,\mathbf{c})}\mathcal{C}\times T_{(y,\mathbf{c})}\mathcal{C}$ satisfies
$D\Xi=\langle\nabla_{x}V(x;\mathbf{c}),\dot{x}\rangle+\langle D_{\mathbf{c}}V(x;\mathbf{c}),\dot{\mathbf{c}}_{1}\rangle-\langle\nabla_{y}V(y;\mathbf{c}),\dot{y}\rangle-\langle D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}_{2}\rangle.$
Because $(x,\mathbf{c})$ and $(y,\mathbf{c})$ lie in $\mathcal{C}$, $\nabla_{x}V=\nabla_{y}V=0$, so the terms with $\dot{x},\dot{y}$ vanish. Moreover, since $\mathbf{c}$ is the same for both points, we have $\dot{\mathbf{c}}_{1}=\dot{\mathbf{c}}_{2}=\dot{\mathbf{c}}$ (the tangent space of $\mathcal{C}^{(2)}$ forces this identification). Thus
$D\Xi=\langle D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}\rangle.$
The vectors $D_{\mathbf{c}}V(x;\mathbf{c})=(e^{2\pi im\cdot x})_{|m|\leq n}$ and $D_{\mathbf{c}}V(y;\mathbf{c})=(e^{2\pi im\cdot y})_{|m|\leq n}$ are linearly independent for $x\neq y$ because the exponentials are linearly independent. Hence $D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c})\neq 0$, and by varying $\dot{\mathbf{c}}$ we obtain a surjective derivative onto $\mathbb{R}$. Therefore $\Xi$ is transverse to $\{0\}$, and $\mathcal{E}$ is a smooth submanifold of $\mathcal{C}^{(2)}$ of codimension 11. Consequently,
$\dim\mathcal{E}=\dim\mathcal{C}^{(2)}-1=2N-d-1.$
For sufficiently high degree $n$, we have $N>d+1$, so $\dim\mathcal{E}<N$. The projection $\pi|_{\mathcal{E}}:\mathcal{E}\to\mathbb{R}^{N}$ is smooth. By Sard’s theorem, the set of critical values of $\pi|_{\mathcal{E}}$ has measure zero in $\mathbb{R}^{N}$. The set of parameters $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ has non-unique extrema is contained in this set of critical values (if $((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{E}$ is a regular point of $\pi|_{\mathcal{E}}$, then nearby parameters have distinct values; non-uniqueness can only occur at critical values). Hence $\mathcal{B}^{(ii)}$ has Lebesgue measure zero.
Conclusion. Both $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ have measure zero, therefore their union also has measure zero. This completes the proof that for almost every $\mathbf{c}$, $V(\cdot;\mathbf{c})$ is a Morse function with unique global extrema. ∎
3.2 Genericity of Cartan Estimates
Let
$V(x;\mathbf{c})=\sum_{|m|\leq n}c_{m}e^{2\pi im\cdot x},\qquad x\in\mathbb{T}^{d},\;\mathbf{c}\in\mathbb{R}^{N},$
where $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$. For a fixed small $\delta>0$ define the complex torus $\mathbb{T}^{d}_{\delta}=\{z\in\mathbb{C}^{d}/\mathbb{Z}^{d}:|\operatorname{Im}z_{j}|\leq\delta\}$ and the analytic norm $\|f\|_{\mathcal{A}}=\sup_{z\in\mathbb{T}^{d}_{\delta}}|f(z)|$. For $\mathbf{c}$ in a compact ball $B_{R}=\{\mathbf{c}:\|\mathbf{c}\|\leq R\}$ the analytic norm of $V(\cdot;\mathbf{c})$ is uniformly bounded.
For condition (iii) we fix $h\in\mathbb{T}^{d}$ and define
$F_{h}(x;\mathbf{c})=V(x+h;\mathbf{c})-V(x;\mathbf{c}),\qquad G_{h}(x;\mathbf{c})=\det\begin{pmatrix}\partial_{x_{i}}V(x)&\partial_{x_{j}}V(x)\\ \partial_{x_{i}}V(x+h)&\partial_{x_{j}}V(x+h)\end{pmatrix}$,
where $(i,j)$ is a fixed pair; the same arguments apply to all finitely many pairs.
Proposition 3.2 (Parameterized Cartan estimate).
For every $R>0$ there exist constants $C_{R}>0$ and $\gamma_{R}>0$ (depending only on $R,d,n$) such that for any $h\in\mathbb{T}^{d}$ and any $\mathbf{c}\in B_{R}$ (except possibly a set of $\mathbf{c}$ where $F_{h}$ or $G_{h}$ is identically zero, which has measure zero), we have
$\operatorname{mes}\bigl\{x\in\mathbb{T}^{d}:\min(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|)<\epsilon\bigr\}\leq C_{R}\,\epsilon^{\gamma_{R}},\qquad\forall\epsilon>0.$ (3.1)
The constants work uniformly for all $h\in\mathbb{T}^{d}$ because the analytic norms of $F_{h},G_{h}$ are bounded uniformly in $h$ and $\mathbf{c}\in B_{R}$.
Proof.
For fixed $h$ and $\mathbf{c}$, the functions $x\mapsto F_{h}(x;\mathbf{c})$, $G_{h}(x;\mathbf{c})$ are real analytic on $\mathbb{T}^{d}$ and extend analytically to $\mathbb{T}^{d}_{\delta}$. Their analytic norms are bounded by $C_{0}(1+\|\mathbf{c}\|)$ for some $C_{0}$, hence uniformly on $B_{R}$. The classical Cartan estimate (see e.g. [8, 13]) gives
$\operatorname{mes}\{x:|f(x)|<\epsilon\}\leq C\,\epsilon^{\gamma}$
for any analytic function $f$ with $\|f\|_{\mathcal{A}}\leq M$, where $C,\gamma>0$ depend only on $M$ and $d$. Applying this to $f=F_{h}$ and $f=G_{h}$ separately yields the estimate for $\min(|F_{h}|,|G_{h}|)$ with constants depending only on the common bound of $\|F_{h}\|_{\mathcal{A}},\|G_{h}\|_{\mathcal{A}}$ over $\mathbf{c}\in B_{R}$. Because $B_{R}$ is compact, these bounds are uniform, and the constants $C_{R},\gamma_{R}$ can be chosen independently of $h$ as well (the family $\{F_{h},G_{h}\}_{h\in\mathbb{T}^{d}}$ is equicontinuous). ∎
For a fixed $h$ and $K\in\mathbb{N}$ define the set of “bad” parameters
$\mathcal{B}^{(iii)}(h,K)=\Bigl\{\mathbf{c}\in\mathbb{R}^{N}:\operatorname{mes}\{x:\min(|F_{h}|,|G_{h}|)<e^{-K}\}>e^{-K^{\mathfrak{c}_{1}}}\Bigr\},$
where $\mathfrak{c}_{1}>0$ is the constant appearing in Definition 2.1(iii) (we assume $\mathfrak{c}_{1}<1$; if $\mathfrak{c}_{1}\geq 1$ one can replace $K^{\mathfrak{c}_{1}}$ by cKcK with $c<1$ by a simple rescaling, which does not affect the genericity statement).
Proposition 3.3 (Integral bound for $\mathcal{B}^{(iii)}(h,K)$).
For each fixed $h\in\mathbb{T}^{d}$ and $R>0$ there exists a constant $C>0$ (depending on $R$) such that for all $K\geq 1$,
$\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)\leq C\,e^{-\gamma_{R}K+K^{\mathfrak{c}_{1}}},$ (3.2)
where $\gamma_{R}>0$ is the exponent from Proposition 3.2. In particular, if $\mathfrak{c}_{1}<1$ then for large $K$ the right-hand side is $\leq Ce^{-\kappa K}$ for some $\kappa>0$.
Proof.
Let $\Psi_{h,K}(\mathbf{c})=\operatorname{mes}\{x\in\mathbb{T}^{d}:\min(|F_{h}(x;\mathbf{c})|,|G_{h}(x;\mathbf{c})|)<e^{-K}\}$. By Proposition 3.2, there exist constants $C_{R},\gamma_{R}>0$ such that for every $\mathbf{c}\in B_{R}$ (excluding a null set $N_{h,R}$ where $F_{h}$ or $G_{h}$ vanishes identically), we have
$\Psi_{h,K}(\mathbf{c})\leq C_{R}e^{-\gamma_{R}K}\qquad\forall K\geq 1.$
The exceptional set $N_{h,R}$ has measure zero and does not affect any integral. Hence
$\int_{B_{R}}\Psi_{h,K}(\mathbf{c})\,d\mathbf{c}\leq C_{R}e^{-\gamma_{R}K}\cdot\operatorname{mes}(B_{R})=:C^{\prime}e^{-\gamma_{R}K},$
where $C^{\prime}=C_{R}\operatorname{mes}(B_{R})<\infty$.
If $\mathbf{c}\in\mathcal{B}^{(iii)}(h,K)\cap B_{R}$, then by definition $\Psi_{h,K}(\mathbf{c})>e^{-K^{\mathfrak{c}_{1}}}$. Consequently,
$e^{-K^{\mathfrak{c}_{1}}}\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)\leq\int_{\mathcal{B}^{(iii)}(h,K)\cap B_{R}}\Psi_{h,K}(\mathbf{c})\,d\mathbf{c}\leq\int_{B_{R}}\Psi_{h,K}(\mathbf{c})\,d\mathbf{c}\leq C^{\prime}e^{-\gamma_{R}K}.$
Rearranging yields
$\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)\leq C^{\prime}e^{-\gamma_{R}K+K^{\mathfrak{c}_{1}}}.$
Taking $C=C^{\prime}$ proves the main inequality. If $\mathfrak{c}_{1}<1$, then for sufficiently large $K$ we have $\gamma_{R}K+K^{\mathfrak{c}_{1}}\leq-\frac{\gamma_{R}}{2}K$. Hence the right-hand side is bounded by $Ce^{-\kappa K}$ with $\kappa=\gamma_{R}/2$. ∎
From now on we assume $\mathfrak{c}_{1}<1$; then there exist constants $\kappa>0$ and $K_{1}$ such that for all $K\geq K_{1}$,
$\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)\leq Ce^{-\kappa K}.$
Proposition 3.4 (Borel-Cantelli for fixed $h$).
For every fixed $h\in\mathbb{T}^{d}$ and $R>0$, the set
$\mathcal{N}_{h,R}=\bigcap_{K_{0}\geq 1}\bigcup_{K\geq K_{0}}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)$
has Lebesgue measure zero. In other words, for almost every $\mathbf{c}\in B_{R}$ there exists $K_{*}(\mathbf{c},h)$ such that $\mathbf{c}\notin\mathcal{B}^{(iii)}(h,K)$ for all $K\geq K_{*}(\mathbf{c},h)$.
Proof.
By Proposition 3.3, there exist constants $C>0$ and $\gamma_{R}>0$ such that for all $K\geq 1$,
$\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)\leq C\,e^{-\gamma_{R}K+K^{\mathfrak{c}_{1}}}.$
Recall that the constant $\mathfrak{c}_{1}$ in Definition 2.1 satisfies $0<\mathfrak{c}_{1}<1$ (this is a standard assumption; see [13]). Hence $K^{\mathfrak{c}_{1}}=o(K)$ as $K\to\infty$, and there exists $\kappa>0$ (e.g., $\kappa=\gamma_{R}/2$) and an integer $K_{1}$ such that for all $K\geq K_{1}$,
$\gamma_{R}K+K^{\mathfrak{c}_{1}}\leq-\kappa K.$
Consequently,
$\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)\leq Ce^{-\kappa K}\qquad\forall K\geq K_{1}.$
Therefore the series
$\sum_{K=1}^{\infty}\operatorname{mes}\bigl(\mathcal{B}^{(iii)}(h,K)\cap B_{R}\bigr)$
converges (the first finitely many terms are finite, and the tail is bounded by a convergent geometric series).
Now apply the first Borel-Cantelli lemma (Proposition 2.7) to the sequence of measurable sets $A_{K}=\mathcal{B}^{(iii)}(h,K)\cap B_{R}$ within the finite-measure space $B_{R}$. The convergence of the sum implies that
$\operatorname{mes}\Bigl(\bigcap_{K_{0}\geq 1}\bigcup_{K\geq K_{0}}A_{K}\Bigr)=0.$
The set $\bigcap_{K_{0}\geq 1}\bigcup_{K\geq K_{0}}A_{K}$ is precisely the set of points that belong to infinitely many $A_{K}$, i.e., $\mathcal{N}_{h,R}$. Hence $\operatorname{mes}(\mathcal{N}_{h,R})=0$. For any $\mathbf{c}\in B_{R}\setminus\mathcal{N}_{h,R}$, there are only finitely many $K$ with $\mathbf{c}\in A_{K}$; taking $K_{*}(\mathbf{c},h)$ larger than the maximal such $K$ yields $\mathbf{c}\notin\mathcal{B}^{(iii)}(h,K)$ for all $K\geq K_{*}(\mathbf{c},h)$. This completes the proof. ∎
Proposition 3.5 (Uniformity in $h$).
There exists a set $\mathcal{Z}\subset\mathbb{R}^{N}$ of Lebesgue measure zero such that for every $\mathbf{c}\notin\mathcal{Z}$ and every $h\in\mathbb{T}^{d}$, condition (iii) is satisfied (i.e. there exists $K_{0}(\mathbf{c},h)$ with the required estimate for all $K\geq K_{0}(\mathbf{c},h)$).
Proof.
Let $\{h_{\ell}\}_{\ell\geq 1}$ be a countable dense subset of $\mathbb{T}^{d}$ (e.g. points with rational coordinates). For each $\ell$ and each integer $m\geq 1$, Proposition 3.4 applied to $R=m$ gives a null set $Z_{\ell,m}\subset B_{m}$ such that for every $\mathbf{c}\in B_{m}\setminus Z_{\ell,m}$ the Cartan estimate holds for $h=h_{\ell}$. Define $Z_{\ell}=\bigcup_{m\geq 1}Z_{\ell,m}$, which is a null set (countable union of null sets). Then for every $\mathbf{c}\notin Z_{\ell}$ (and hence for every $\mathbf{c}\in\mathbb{R}^{N}\setminus Z_{\ell}$) the estimate holds for $h=h_{\ell}$. Set $\mathcal{Z}=\bigcup_{\ell\geq 1}Z_{\ell}$; then $\mathcal{Z}$ is null.
Now fix $\mathbf{c}\notin\mathcal{Z}$ and an arbitrary $h\in\mathbb{T}^{d}$. Choose a sequence $h_{\ell_{k}}\to h$. Because $F_{h}$ and $G_{h}$ depend continuously on $h$ (uniformly in $x$ on the compact $\mathbb{T}^{d}$), for any $\epsilon>0$ we can find $k$ large enough so that
$\|F_{h}-F_{h_{\ell_{k}}}\|_{\infty}\leq\epsilon,\qquad\|G_{h}-G_{h_{\ell_{k}}}\|_{\infty}\leq\epsilon.$
Then for any $K$,
$\min(|F_{h}|,|G_{h}|)\leq\min(|F_{h_{\ell_{k}}}|,|G_{h_{\ell_{k}}}|)+\epsilon.$
Hence if $\min(|F_{h_{\ell_{k}}}|,|G_{h_{\ell_{k}}}|)\geq 2\epsilon$, then $\min(|F_{h}|,|G_{h}|)\geq\epsilon$. Taking $\epsilon=e^{-K}$ we obtain that the set where $\min(|F_{h}|,|G_{h}|)<e^{-K}$ is contained in the set where $\min(|F_{h_{\ell_{k}}}|,|G_{h_{\ell_{k}}}|)<2e^{-K}$. Consequently,
$\operatorname{mes}\{x:\min(|F_{h}|,|G_{h}|)<e^{-K}\}\leq\operatorname{mes}\{x:\min(|F_{h_{\ell_{k}}}|,|G_{h_{\ell_{k}}}|)<2e^{-K}\}.$
Since $\mathbf{c}\notin\mathcal{Z}$, it belongs to the good set for $h_{\ell_{k}}$. Hence there exists $K_{*}(\mathbf{c},h_{\ell_{k}})$ such that for all $K\geq K_{*}$,
$\operatorname{mes}\{x:\min(|F_{h_{\ell_{k}}}|,|G_{h_{\ell_{k}}}|)<2e^{-K}\}\leq C_{R}(2e^{-K})^{\gamma_{R}},$
where $R$ is chosen so that $\mathbf{c}\in B_{R}$ (e.g., $R=\|\mathbf{c}\|+1$), and $C_{R},\gamma_{R}$ are the constants from Proposition 3.2. Because $\mathfrak{c}_{1}<1$ (see Definition 2.1), the right-hand side is $\leq e^{-K^{\mathfrak{c}_{1}}}$ for all sufficiently large $K$ (say $K\geq K_{1}$). Taking $K_{0}(\mathbf{c},h)=\max(K_{*},K_{1})$ gives the desired estimate for $h$. Thus condition (iii) holds for every $h\in\mathbb{T}^{d}$. ∎
The same reasoning applies to condition (iv). Define
$\mathcal{B}^{(iv)}(\eta,h_{0},K)=\Bigl\{\mathbf{c}:\operatorname{mes}\{x:\min(|V(x;\mathbf{c})-\eta|,|\langle\nabla V(x;\mathbf{c}),h_{0}\rangle|)<e^{-K}\}>e^{-K^{\mathfrak{c}_{1}}}\Bigr\}.$
For fixed $R$ and compact intervals $I_{R}$ containing the range of $V$ on $B_{R}$, the functions $F_{\eta}=V-\eta$ and $G_{h_{0}}=\langle\nabla V,h_{0}\rangle$ are analytic in $x$ and linear in $\mathbf{c},\eta$ (for $\eta$ in $I_{R}$). The same Fubini-Borel-Cantelli argument yields that for each $(\eta,h_{0})$ there is a null set outside which the gradient Cartan estimate holds. Taking a countable dense set $\{(\eta_{p},h_{0,q})\}$ in $I_{R}\times S^{d-1}$ and using the continuity in $(\eta,h_{0})$ (uniform on compact sets) we obtain a null set $\mathcal{Z}^{\prime}$ such that for every $\mathbf{c}\notin\mathcal{Z}^{\prime}$ the gradient estimate holds for all $(\eta,h_{0})$. Details are completely analogous to Proposition 3.5 and are omitted.
Lemma 3.6 (Genericity of Cartan Estimates).
The set of coefficients $\mathbf{c}\in\mathbb{R}^{N}$ for which the potential $V(\cdot;\mathbf{c})$ fails to satisfy the Cartan estimates (conditions (iii) and (iv) of Definition 2.1) has Lebesgue measure zero.
Proof.
We first treat condition (iii). For each $R>0$, let $\mathcal{Z}_{R}$ be the null set obtained from Proposition 3.5 (which depends on $R$ because the constants $C_{R},\gamma_{R}$ in Proposition 3.2 depend on $R$). Actually Proposition 3.5 was proved for a fixed $R$, so we denote by $\mathcal{Z}_{R}$ the null set such that for every $\mathbf{c}\notin\mathcal{Z}_{R}$ and every $h\in\mathbb{T}^{d}$, condition (iii) holds. Set $\mathcal{Z}^{(iii)}=\bigcup_{R\in\mathbb{N}}\mathcal{Z}_{R}$; then $\mathcal{Z}^{(iii)}$ is a countable union of null sets, hence null. For any $\mathbf{c}\notin\mathcal{Z}^{(iii)}$, there exists $R$ with $\|\mathbf{c}\|\leq R$, so $\mathbf{c}\notin\mathcal{Z}_{R}$ and condition (iii) holds for all $h$.
Similarly, for condition (iv) we obtain a null set $\mathcal{Z}^{(iv)}$ (by taking the union over $R$ of the null sets from the gradient version of Proposition 3.5). Finally, let $\mathcal{Z}=\mathcal{Z}^{(iii)}\cup\mathcal{Z}^{(iv)}$. Then $\mathcal{Z}$ is null, and for every $\mathbf{c}\notin\mathcal{Z}$ both Cartan estimates (iii) and (iv) are satisfied. Therefore the set of coefficients for which the estimates fail is contained in $\mathcal{Z}$ and has Lebesgue measure zero. ∎
Proof of Theorem $2.$
Recall that $\mathfrak{G}$ is defined by four conditions (Definition 2.1). Let $\mathcal{B}=\mathcal{B}^{(i)}\cup\mathcal{B}^{(ii)}\cup\mathcal{B}^{(iii)}\cup\mathcal{B}^{(iv)}$ be the set of parameters where at least one condition fails. By Lemma 3.1, $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ have measure zero. By Lemma 3.6, $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ also have measure zero. A finite union of measure-zero sets has measure zero, so $\mathcal{B}$ is null. Hence the set of good parameters $\mathcal{G}=\mathbb{R}^{N}\setminus\mathcal{B}$ has full Lebesgue measure. This proves the genericity conjecture of Goldstein, Schlag, and Voda [13]. ∎
Remark 3.7 (Topological Genericity).
The proof of Theorem 2.2 actually establishes a stronger result: the set of good parameters $\mathcal{G}$ is not only of full measure but also a residual set in $\mathbb{R}^{N}$. This follows because the exceptional sets $\mathcal{B}^{(i)}$ and $\mathcal{B}^{(ii)}$ are contained in smooth submanifolds of positive codimension, which are closed and have empty interior (hence nowhere dense). The exceptional sets $\mathcal{B}^{(iii)}$ and $\mathcal{B}^{(iv)}$ can also be shown to be of first category. Thus $\mathcal{G}$ is the complement of a finite union of first category sets, making it residual. This confirms that the potentials in class $\mathfrak{G}$ are generic in both the measure-theoretic and topological senses.
Acknowledgments
This work was supported by NSFC (No. 11571327, 11971059).